# CLEF 2024: Conference and Labs of the Evaluation Forum, September 09–12, 2024, Grenoble, France

### v1

@@ 1 Introduction
The effectiveness of LLMs in information extraction (IE) tasks is deeply influenced by the context provided during the input phase. In this context, “context” refers to the segments of the scholarly articles from which the (Task, Dataset, Metric, Score) data is extracted. Unlike the concept of in-context learning, where models are primed with examples, our focus is on selecting relevant sections of text to guide the LLMs towards accurate and relevant information extraction while minimizing the risk of generating irrelevant or false information (hallucinations).
@@
Longer contextual inputs can lead to decreased model accuracy and higher rates of hallucination due to the potential distraction of irrelevant information [3, 4]. Therefore, identifying the optimal length and specificity of context is crucial [3]. This study explores three distinct types of contextual inputs to balance these factors effectively: DocTAET (Document Title, Abstract, Experimental Setup, and Tabular Information), DocREC (Results, Experiments, and Conclusions), and DocFULL (the entire document).
To contribute to this effort, we participate in the “SOTA? Tracking the State-of-the-Art in Scholarly Publications” shared task as Task 4 in the SimpleText [^1] https://simpletext-project.com/2024/en/ track of CLEF 2024. The goal of the SOTA? shared task is to develop systems that, given the full text of an AI paper, can recognize whether the paper reports model scores on benchmark datasets and, if so, extract all pertinent (Task, Dataset, Metric, Score) tuples presented within the paper.
@@ 2 Related Work
In the realm of information extraction, the application of large language models such as GPT-3.5 and LLAMA has demonstrated significant promise. A study by Hu et al. (2024) highlighted the capabilities of GPT-3.5 and GPT-4 in clinical named entity recognition (NER), showcasing their adeptness at processing intricate clinical datasets with limited prerequisite training. Through strategic prompt engineering, these models exhibited remarkable improvements in performance for extracting medical entities from clinical documentation, reinforcing the potential of LLMs in executing complex NER tasks within the healthcare sector [13].
The emergence of large language models, including ChatGPT, has sparked a reconsideration of specialized versus general-purpose training approaches in the context of LLMs. Building upon the foundation of utilizing large language models for specialized information extraction tasks, "Large Language Models for Scientific Information Extraction: An Empirical Study for Virology" delves into the utilization of these models in the domain of virology. The study showcases how LLMs, specifically tuned for scientific content, can efficiently parse and extract virology-related information from a plethora of scientific publications. This research underlines the importance of fine-tuning and prompt engineering in enhancing the model’s ability to discern relevant scientific facts, contributing to the development of domain-specific leaderboards. The approach exemplified in this paper demonstrates an effective strategy for context selection in LLM-based leaderboard generation, emphasizing the necessity for domain-specific adjustments to maximize the accuracy and relevance of the extracted information.
Our work, "Effective Context Selection in LLM-Based Leaderboard Generation: An Empirical Study," extends these findings by empirically examining various context selection strategies in the generation of leaderboards across multiple scientific domains. We analyze the impact of tailored context cues and the integration of domain-specific knowledge on the precision and utility of automatically generated leaderboards. Our contributions include a comprehensive evaluation framework for assessing the effectiveness of different context selection strategies and the development of a novel methodology that significantly enhances the accuracy and relevance of LLM-based leaderboard generation. This study not only corroborates the pivotal role of context in leveraging LLMs for information extraction but also introduces innovative techniques that refine the process of generating insightful and reliable leaderboards in the scientific community.
@@ 3 Methodology
This section details the methodology employed in our study, encompassing data collection, preprocessing, model selection, and evaluation metrics. The goal is to systematically evaluate the performance of three state-of-the-art LLMs—Mistral 7B, Llama-2, GPT-4-Turbo and GPT-4.o—across different context types for the task of leaderboard extraction.
@@ 3.1 Data Collection and Preprocessing
We utilized data that was previously published by [NLDB]. The dataset consists of (T, D, M, S) annotations for thousands of AI articles available on PwC (CC BY-SA). These articles cover various AI domains such as Natural Language Processing, Computer Vision, Robotics, Graphs, Reasoning, etc., making them representative for empirical AI research. The specific PwC source was downloaded on December 09, 2023. The corpus comprised over 8,000 articles, with 7,987 used for training and 994 for testing. These articles, originally sourced from arXiv under CC-BY licenses, are available as latex code source, each accompanied by one or more (T, D, M, S) annotations from PwC. The articles’ metadata was directly obtained from the PwC data release, and the articles collection had to be reconstructed by downloading them from arXiv under CC-BY licenses.
@@
In this work, for the first time, we empirically investigate the impact of leveraging different contexts to the LLM. Specifically, we test the following three main types:
@@ 3.2 DocTAET
Delineated in prior work [11], the context to the LLM comprises text selected from the (T)-title, (A)-abstract, (E)-experimental setup, and (T)-tabular information parts of the full-text. It yields an average context length of 493 words, ranging from a minimum of 26 words to a maximum of 7,361 words. These specific selections targeted the areas of the paper where the (T, D, M, S) were most likely to be found.
@@ 3.3 DocREC
Introduced for the first time in this work, the DocREC context comprises text selected from the sections named (R)-results, (E)-experiments, and (C)-conclusions with allowances for variations in the three names. Complementary but still unique to DocTAET, the DocREC context representation aims to distill the essence of the research findings and conclusions into a succinct format. This context, ended by being much longer than DocTAET, yielded an average length of 1,586 words, with a minimum length of 27 words and a maximum length of 127,689 words
@@ 3.5 Models
In this work, we selected two open-sourced models as well as a closed-source model.
Mistral-7B. As the first open-sourced model, we selected Mistral-7B [mistral]. This model as the name suggested is a 7-billion-parameter language model optimized for performance and efficiency. It introduces the Grouped-Query Attention (GQA) for rapid inference and reduced memory requirements, and Sliding Window Attention (SWA) for handling long sequences with lower computational costs. The model surpasses existing models in benchmarks, including reasoning, mathematics, and code generation tasks. It also features fine-tuning capabilities for instruction following, achieving superior performance in human and automated benchmarks. Mistral 7B is designed for real-time applications, supports high throughput, and its architecture enables effective sequence generation with optimized memory usage. The model is released under the Apache 2.0 license, with its source code on Github, facilitating broad accessibility and application in various tasks.
@@ 4 Evaluations
Experimental setup. For training, we had one main experimental setting based on the 15 instructions. As elicited earlier in subsection 3.1, each of the 15 instruction were instantiated with the 7,987 (T, D, M, S) data instances and the SOTA question resulting in a total of 119,805 instances to instruction finetune Flan-T5 Large. In this scenario, we hypothesized that this repetition in the data instances across the instructions would cause the resulting model to overfit the training dataset. Thus to control for this, we applied the following experimental setup. Each instruction was instantiated with a random selection of only half the 7,987 (T, D, M, S) data instances resulting in a finetuning dataset of a sizeable 41,340 instances that had leaderboards. The papers w/o leaderboards were also similarly halved. In the test scenario, however, we report per instruction (T, D, M, S) instantiated data results. As shown in Table 1, for a few-shot test set had 994 (T, D, M, S) instances and 648 papers without leaderboards and zero-shot test set had 652 (T, D, M, S) instances and 548 papers without leaderboards, evaluations results are shown for each instruction separately with a total of 2,353 underlying papers representing those with and without leaderboards. Model hyperparameter details are in Appendix C. In terms of compute, all experiments including inference were run on an NVIDIA 3090 GPU. Training took  20 hours on the 50% sampled dataset, while inference lasted  10 minutes for 2,353 test instances.
@@
Table 2: Evaluation results of Llama-2, Mistral, GPT-4-Turbo and GPT-4.o w.r.t with output evaluations as a structured summary generation task (reported with ROUGE metrics) as well as binary classification between papers with and without leaderboards (reported as General Accuracy) for each of the 15 instructions from DROP and SQuAD datasets, respectively, with 50% randomly selected paper instantiations per template.
@@
Table 3: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.o w.r.t. the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output in terms of F1 score for each of the 15 instructions from DROP and SQuAD datasets, respectively, with 50% randomly selected paper instantiations per template.
@@
Table 4: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.o w.r.t. the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output in terms of Precision score for each of the 15 instructions from DROP and SQuAD datasets, respectively, with 50% randomly selected paper instantiations per template.
@@ 4.1 Results and Discussion
    Few-shot Precision (Table 3):
    • Llama-2 7B: Shows a balanced performance between exact and partial metrics. For example, it achieves an Overall score of 26.4 in the partial mode under DocTAET context, indicating its ability to handle partial matches effectively.
    • Mistral 7B: Excels in both exact and partial metrics, with the highest Overall score of 28.88 in the partial mode under DocTAET context. This model provides the best trade-off between precision and other performance metrics, especially in few-shot settings.
    • GPT-4-Turbo: Shows moderate performance with an Overall score of 21.65 in the partial mode under DocFULL context. It requires further optimization to compete with other models.
@@
    Zero-shot Precision (Table 3):
@@
    • Mistral 7B: Continues to lead with the highest Overall score of 28.33 in the partial mode under DocTAET context, reinforcing its capability in zero-shot scenarios.
    • GPT-4-Turbo: Provides moderate performance in partial mode with an Overall score of 15.94 under DocFULL context, highlighting areas for improvement.
@@
    • Llama-2 7B: Achieves high precision scores, particularly in the partial mode with an Overall score of 32.32 in few-shot settings under DocTAET context.
@@
    • GPT-4-Turbo: Shows competitive precision scores in few-shot settings, particularly in the partial mode with an Overall score of 30.69 under DocFULL context, but requires improvement in zero-shot settings.
    • GPT-4.o: Shows competitive precision scores in few-shot settings, particularly in the partial mode with an Overall score of 34.90 under DocTAET context.
@@ 5 Conclusions
Our findings highlight that fine-tuning open-source models like Mistral 7B and Llama-2 7B can yield competitive, and in some cases superior, results compared to the proprietary model GPT-4.o and GPT-4-Turbo. The DocTAET context typically yields dependable and accurate performance, while the DocREC context excels in scenarios where precision is paramount. The interaction among different context types and their consequent impact on model performance offers extensive insights for ongoing research and the practical deployment of language models.
The ramifications of our research are substantial, indicating that with meticulous context design and implementation, fine-tuned open-source LLMs are well-equipped for tracking and synthesizing scientific progress, thus providing a current and nuanced leaderboard for any given academic field. The potential of this technology to support and augment the endeavors of academics and policymakers is significant, heralding novel opportunities for automated, real-time meta-analysis.
While this research makes significant progress in elucidating and enhancing context selection for LLMs, it also unveils several avenues for future exploration. Delving into hybrid context selection methods, domain-specific adaptations, and the assimilation of structured data could lead to even more sophisticated leader- board generation. As LLMs continue to advance, the pursuit of optimizing their potential remains a dynamic and impactful field of research.

### v2

@@ 1 Introduction
The effectiveness of LLMs in information extraction (IE) tasks is deeply influenced by the context provided during the input phase. In this context, “context” refers to the segments of the scholarly articles from which the (Task, Dataset, Metric, Score) data is extracted. Unlike the concept of in-context learning, where models are primed with examples, our focus is on selecting relevant sections of text to guide the LLMs toward accurate and relevant information extraction while minimizing the risk of generating irrelevant or false information (hallucinations).
@@
Existing research has shown that longer contextual inputs can lead to decreased model accuracy and higher rates of hallucination due to the potential distraction of irrelevant information [3, 4]. Therefore, identifying the optimal length and specificity of context is crucial [3]. This study explores three distinct types of contextual inputs to empirically evaluate their impact on the task of extracting leaderboards: DocTAET (Document Title, Abstract, Experimental Setup, and Tabular Information), DocREC (Results, Experiments, and Conclusions), and DocFULL (the entire document).
@@
Additionally, we participate in the “SOTA? Tracking the State-of-the-Art in Scholarly Publications” shared Task 4 in the SimpleText [^1] https://simpletext-project.com/2024/en/ track of CLEF 2024 [8, 9]. The goal of the SOTA? shared task is to develop systems that, given the full text of an AI paper, can recognize whether the paper reports model scores on benchmark datasets and, if so, extract all pertinent (Task, Dataset, Metric, Score) tuples presented within the paper.
@@ 2 Related Work
In the realm of information extraction, the application of LLMs such as GPT-3.5 and LLAMA has demonstrated significant promise. A study by [15] highlighted the capabilities of GPT-3.5 and GPT-4 in clinical named entity recognition (NER), showcasing their adeptness at processing intricate clinical datasets with limited prerequisite training. Through strategic prompt engineering, these models exhibited remarkable improvements in performance for extracting medical entities from clinical documentation, reinforcing the potential of LLMs in executing complex NER tasks within the healthcare sector [15].
The emergence of LLMs, including ChatGPT, has sparked a reconsideration of specialized versus general-purpose training approaches in the context of LLMs. Building upon the foundation of utilizing LLMs for specialized information extraction tasks, [16] delves into the utilization of these models in the domain of virology. The study showcases how LLMs, specifically tuned for scientific content, can efficiently parse and extract virology-related information from a plethora of scientific publications. This research underlines the importance of fine-tuning and prompt engineering in enhancing the model’s ability to discern relevant scientific facts, contributing to the development of domain-specific leaderboards. The approach exemplified in this paper demonstrates an effective strategy for context selection in LLM-based leaderboard generation, emphasizing the necessity for domain-specific adjustments to maximize the accuracy and relevance of the extracted information.
In our previous study [17], we expanded on these findings by empirically investigating different ways to select context in creating leaderboards. We analyze the impact of tailored context cues and the integration of domain-specific knowledge on the precision and utility of automatically generated leaderboards. Our contributions included a thorough assessment of the latest LLMs, both open-source and proprietary, for assessing the effectiveness of different context selection strategies and the development of a novel methodology that significantly enhances the accuracy and relevance of LLM-based leaderboard generation. This study not only corroborates the pivotal role of context in leveraging LLMs for information extraction but also introduces innovative techniques that refine the process of generating insightful and reliable leaderboards in the scientific community.
@@ 3 Methodology
This section details the methodology employed in our study, encompassing data collection, preprocessing, model selection, and evaluation metrics. The goal is to systematically evaluate the performance of four state-of-the-art LLMs—Mistral 7B, Llama-2, GPT-4-Turbo and GPT-4.o—across different context types for the task of leaderboard extraction.
@@ 3.1 Data Collection and Preprocessing
We utilized data that was previously published by [17] and re-released as the SOTA? shared task training corpus. The dataset consists of (T, D, M, S) annotations for thousands of AI articles available on PwC (CC BY-SA). These articles cover various AI domains such as Natural Language Processing, Computer Vision, Robotics, Graphs, Reasoning, etc., making them representative for empirical AI research. The specific PwC source was downloaded on December 09, 2023. The corpus comprised over 8,000 articles, with 7,987 used for training and 994 for testing, including 751 in the few-shot setting and 241 in the zero-shot setting. These articles, originally sourced from arXiv under CC-BY licenses, are available as latex source code, each accompanied by one or more (T, D, M, S) annotations from PwC. The articles’ metadata was directly obtained from the PwC data release, and the articles collection had to be reconstructed by downloading them from arXiv under CC-BY licenses.
@@
As introduced in our prior work [17], we compared the performances of the four state-of-the-art LLMs on the following three contexts:
@@ 3.2 DocTAET
Delineated in prior work [13], the context to the LLM comprises text selected from the (T)-title, (A)-abstract, (E)-experimental setup, and (T)-tabular information parts of the full-text. It yields an average context length of 493 words, ranging from a minimum of 26 words to a maximum of 7,361 words. These specific selections targeted the areas of the paper where the (T, D, M, S) were most likely to be found. An example of this context selection is illustrated in  Figure 2.
@@ 3.3 DocREC
Introduced for the first time in this work, the DocREC context comprises text selected from the sections named (R)-results, (E)-experiments, and (C)-conclusions with allowances for variations in the three names. Complementary but still unique to DocTAET, the DocREC context representation aims to distill the essence of the research findings and conclusions into a succinct format. This context, ended by being much longer than DocTAET, yielded an average length of 1,586 words, with a minimum length of 27 words and a maximum length of 127,689 words. An example of this context selection is illustrated in  Figure 3.
@@ 3.5 Models
In this work, we selected two open-sourced models as well as two closed-source models.
Mistral-7B. As the first open-sourced model, we selected Mistral-7B [6]. This model as the name suggested is a 7-billion-parameter language model optimized for performance and efficiency. It introduces the Grouped-Query Attention (GQA) for rapid inference and reduced memory requirements, and Sliding Window Attention (SWA) for handling long sequences with lower computational costs. The model surpasses existing models in benchmarks, including reasoning, mathematics, and code generation tasks. It also features fine-tuning capabilities for instruction following, achieving superior performance in human and automated benchmarks. Mistral 7B is designed for real-time applications, supports high throughput, and its architecture enables effective sequence generation with optimized memory usage. The model is released under the Apache 2.0 license, with its source code on Github, facilitating broad accessibility and application in various tasks.
@@ 4 Evaluations
Experimental setup. For training, we had one main experimental setting based on the 15 instructions. As elicited earlier in subsection 3.1, each of the 15 instruction were instantiated with the 7,987 (T, D, M, S) data instances and the SOTA question resulting in a total of 119,805 instances to instruction finetune out LLMs. In this scenario, we hypothesized that this repetition in the data instances across the instructions would cause the resulting model to overfit the training dataset. Thus to control for this, we applied the following experimental setup. Each instruction was instantiated with a random selection of only half the 7,987 (T, D, M, S) data instances resulting in a finetuning dataset of a sizeable 41,340 instances that had leaderboards. The papers w/o leaderboards were also similarly halved. As shown in Table 1, the few-shot test included 1,401 paper instances, comprising 648 papers without leaderboards and 735 papers with leaderboards. The zero-shot test set consisted of 789 papers, with 548 papers without leaderboards and 241 papers with leaderboards. Model hyperparameter details are in Appendix C. In terms of compute, all experiments including inference were run on an NVIDIA 3090 GPU. Training took  20 hours on the 50% sampled dataset, while inference lasted 10 minutes for  2k test instances.
@@
Table 2: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.0 for the shared task, reported using the metrics proposed for the task. The output evaluations are conducted as a structured summary generation task (reported with ROUGE metrics) and as a binary classification task to distinguish between papers with and without leaderboards (reported as General Accuracy).
@@
Table 3: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.0 for the shared task, reported using the metrics proposed for the task. The evaluation considers the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output, reported in terms of F1 score.
@@
Table 4: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.o w.r.t. the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output in terms of Precision score.
Few-shot Zero-shot Model Mode Task Dataset Metric Score Overall Task Dataset Metric Score Overall Context   Llama-2 7B Exact 15.09 9.42 10.07 2.19 9.19 8.94 4.36 7.51 1.66 5.62 DocREC Partial 2.63 16.23 15.86 2.49 14.30 15.40 10.60 11.70 1.97 9.92 Exact 28.52 16.11 19.32 1.10 16.26 19.52 10.10 15.24 0.69 11.39 DocTAET Partial 41.89 29.33 29.43 1.33 25.50 34.53 20.73 24.31 0.85 20.10 Exact 0.81 0.69 0.48 0.12 0.53 1.06 0.67 0.78 0.17 0.67 DocFULL Partial 1.17 0.93 0.86 0.19 0.79 1.74 1.29 1.29 0.28 1.15   Mistral 7B Exact 20.77 12.14 14.49 4.93 13.08 12.03 7.89 10.38 3.37 8.42 DocREC Partial 30.84 21.12 22.07 5.49 19.88 19.98 14.07 15.45 3.84 13.33 Exact 28.91 16.03 20.99 1.62 16.89 20.79 11.03 16.97 0.93 12.43 DocTAET Partial 40.15 28.36 29.59 1.95 25.01 34.58 21.02 24.83 1.08 20.38 Exact 0.41 0.29 0.29 0.29 0.32 0.11 0.17 0.17 0.39 0.21 DocFULL Partial 0.6 0.43 0.41 0.43 0.47 0.28 0.34 0.39 0.44 0.36   GPT-4-Turbo Exact 4.35 3.53 2.81 2.42 3.28 2.37 2.98 2.15 1.83 2.33 DocREC Partial 9.42 7.98 6.30 4.12 6.95 7.67 6.18 5.68 3.12 5.66 Exact 1.56 1.41 0.50 0.39 0.96 0.58 0.40 0.17 0.06 0.30 DocTAET Partial 3.26 2.84 1.59 0.86 2.14 1.39 0.81 0.81 0.06 0.76 Exact 1.8 1.68 1.05 1.32 1.46 1.29 1.57 0.95 1.46 1.32 DocFULL Partial 3.73 3.40 2.63 2.23 3.00 3.42 3.08 2.35 1.89 2.68   GPT-4.o Exact 14.63 14.60 14.05 9.75 13.26 13.41 12.58 14.57 8.68 12.31 DocREC Partial 34.81 29.58 26.60 14.41 26.35 31.13 26.05 25.06 12.71 23.74 Exact 10.97 9.93 7.72 1.64 7.57 9.47 7.04 7.56 1.62 6.42 DocTAET Partial 24.79 20.75 16.21 2.89 16.16 20.15 16.40 13.39 2.33 13.07 Exact 14.82 12.88 13.00 9.97 12.66 11.42 13.05 14.05 9.76 12.07 DocFULL Partial 32.41 27.48 24.46 15.13 24.87 31.08 24.69 23.91 13.32 23.25
Table 5: Evaluation results of Llama-2, Mistral, GPT-4-Turbo, and GPT-4.o w.r.t. the individual (Task, Dataset, Metric, Score) elements and Overall in the model JSON generated output in terms of Recall score.
@@ 4.1 Results and Discussion
    Few-shot F1 Score analysis (Table 3):
    • Llama-2 7B: Shows a balanced performance between exact and partial metrics. For example, it achieves an Overall score of 26.40 in the partial mode under DocTAET context, indicating its ability to handle partial matches effectively.
    • Mistral 7B: Excels in both exact and partial metrics, with the highest Overall score of 28.88 in the partial mode under DocTAET context. This model provides the best trade-off between precision and other performance metrics.
    • GPT-4-Turbo: Shows moderate performance with an Overall score of 12.13 in the partial mode under DocREC context. It requires further optimization to compete with other models.
@@
    Zero-shot F1 Score analysis (Table 3):
@@
    • Mistral 7B: Continues to lead with the highest Overall score of 16.14 in the exact mode under DocTAET context, reinforcing its capability in zero-shot scenarios.
    • GPT-4-Turbo: Provides moderate performance in partial mode with an Overall score of 10.13 under DocFULL context, highlighting areas for improvement.
@@
    • Llama-2 7B: Achieves moderate precision scores, particularly in the partial mode with an Overall score of 35.56 in zero-shot settings under DocREC context.
@@
    • GPT-4-Turbo: Shows competitive precision scores in zero-shot settings, particularly in the partial mode with an Overall score of 50.20 under DocFULL context.
    • GPT-4.o: Achieves moderate precision scores in few-shot settings, particularly in the exact mode, with an Overall score of 16.40 under DocFULL context.
  4. 4.
    Recall Scores (Table 5):
    • Llama-2 7B: Demonstrates notable recall scores in few-shot settings, particularly in the partial mode with an Overall score of 25.50 under DocTAET context. However, its performance varies significantly across different contexts and modes.
    • Mistral 7B: Leads in recall scores, achieving the highest Overall scores in both exact and partial modes across various contexts. Notably, it achieves an Overall score of 16.89 in the exact mode in few-shot setting and 12.43 in few-shot under DocTAET context.
    • GPT-4-Turbo: Shows moderate recall performance, with room for improvement, particularly in the zero-shot settings. Its best recall score is in the exact mode under DocREC context with an Overall score of 3.28.
    • GPT-4.o: Excels in recall scores, especially in few-shot settings, achieving an Overall score of 26.35 in the partial mode under DocREC context. It maintains competitive recall scores across different contexts and modes.
@@ 5 Conclusions
Our participation in the shared task has demonstrated that fine-tuning open-source models like Mistral 7B and Llama-2 7B can yield competitive, and in some cases superior, results compared to proprietary models such as GPT-4.o and GPT-4-Turbo. Throughout our experiments, the DocTAET context typically delivered dependable and accurate performance, while the DocREC context excelled in scenarios where precision is paramount. The interaction among different context types and their consequent impact on model performance provided valuable insights into ongoing research and practical deployment of LLMs.
The implications of our findings within the context of the shared task are substantial. They indicate that with meticulous context design and implementation, fine-tuned open-source LLMs are well-suited for tracking and synthesizing scientific progress. This capability enables the provision of current and nuanced leaderboards for any given academic field. The potential of this technology to support and augment the efforts of academics and policymakers is significant, heralding novel opportunities for automated or semi-automated leaderboard construction.
While we have made significant progress in elucidating and enhancing context selection for LLMs, several avenues for future exploration have emerged. Investigating hybrid context selection methods, domain-specific adaptations, and the integration of structured data could lead to even more sophisticated leaderboard generation. As LLMs continue to advance, optimizing their potential remains a dynamic and impactful field of research.
In conclusion, our involvement in the shared task has not only highlighted the effectiveness of fine-tuned open-source models but also emphasized the importance of strategic context selection in maximizing model performance. These insights contribute to the broader understanding of LLM capabilities and pave the way for future advancements in the field.
Acknowledgements
This work was jointly supported by the German BMBF project SCINEXT (01lS22070) and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number: NFDI4DataScience (460234259).